OpenStudy (alprincenofl):

Choices: If u=〈1,2〉, v=〈1,0〉 and w=〈-1,2〉, then 〈u+v,w〉= The choices are: 1) 0 2) 2 3) 〈-1,4〉

geerky42 (geerky42):

What is "\(\small\square\)" in "〈u+v,w〉" supposed to be?

OpenStudy (anonymous):

〈u+v,w〉= u+v is vector <2,2> I dont know what the comma is supposed to be, do they want a matrix with columns u+v and w?

OpenStudy (alprincenofl):

@optiquest, Yes it's Martix

OpenStudy (alprincenofl):

@geerky42, There isn't any symbols in the equation

OpenStudy (alprincenofl):

The question is: If u=〈1,2〉, v=〈1,0〉 and w=〈-1,2〉, then 〈u+v,w〉= The choices are: 1) 0 2) 2 3) 〈-1,4〉

OpenStudy (alprincenofl):

The answer is the second one 2 But I don't know how

OpenStudy (alprincenofl):

http://i.imgur.com/reKUmn5.png

OpenStudy (alprincenofl):

@geerky42

OpenStudy (anonymous):

I think this has to do with inner product spaces actually